# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ Appendix E Open-Source Software and Data Availability
  • Medical Ultrasound Tomography. In Refs. [66, 220], the authors develop a three-dimensional anatomically realistic numerical phantom of the female breast and a high-fidelity three-dimensional imaging model of ring-array USCT breast ultrasound. The goal of the generated numerical breast phantoms (NBPs) and corresponding simulated measurements is to provide a dataset to 1) support the development and assessment (using objective measurement of image quality) of model-based and learned image reconstruction methods and 2) conduct virtual imaging trials to guide the design of new USCT breast imaging systems. Compared to other acoustic NBPs in the literature, the ones in Ref. [66] are particularly relevant to the development of new learning-based USCT image reconstruction method since they 1) comprise realistic structures and acoustic properties; 2) include lesions and/or other pathologies; and 3) are representative of the stochastic variability in breast size, shape, composition, anatomy, and tissue properties observed in a specified cohort of to-be-imaged subjects. Leveraging tools developed and released by the Virtual Imaging Clinical Trials for Regulatory Evaluation (VICTRE) project [221, 222] of the Food and Drug Administration (FDA), the NBPs in Ref. [66] correspond to the four different levels of breast density defined according to the American College of Radiology’s (ACR) Breast Imaging Reporting and Data System (BI-RADS) [223]: A) Breast is almost entirely fat, B) Breast has scattered areas of fibroglandular density, C) Breast is heterogeneously dense, and D) Breast is extremely dense. Each NBP is a 3D voxelized map consisting of ten tissue types: fat, skin, glandular, nipple, ligament (connective tissue), muscle, terminal duct lobular unit, duct, artery, and vein. In the framework of the authors, large ensembles of stochastic NBPs with realistic variability in breast volume, shape, fraction of glandular tissue, ligament orientation, tissue anatomy, can be generated by controlling input parameters and selecting the random seed number. A variety of lesions (e.g., circumscribed or spiculated masses) can also be inserted at physiologically plausible locations. Tissue-specific acoustic properties (density, SOS, and AA) are then stochastically assigned to each tissue structure within physiological ranges and acoustic heterogeneity within fatty and glandular tissues is also modeled as spatially correlated random fields. Realization of these numerical phantoms and corresponding simulated measurements can be downloaded from Refs. [224, 225, 226], and additional NBPs can be generated using the open-source software in Ref. [227].
TABLE V: Paper Summary
@@
Reference Pure ML Methods Hybrid Methods Supervised Semi-Supervised Self-Supervised Unsupervised Simulations/ Measurements-Driven Physics Aware Parameterized Solution Plug & Play Seismic Wu and Lin [41] Araya-Polo et al. [1] Renan et al. [47] Wang et al. [228] Kazei et al. [229] Wu et al. [230] Feng et al. [49] Hu et al. [231] Zhu et al. [56] Wu et al. [232] Fu et al. [92] Vamaraju et al. [233] Izzatullah et al. [93] Medical Ultrasound Lozenski et al. [9] Vedula et al. [234] Feigin et al. [235] Zhao et al. [236] Tong et al. [237] Liu et al. [238] Zhang et al. [239] Zhang et al. [240] Dai et al. [241] Dai et al. [242] Fan et al. [243] Acoustic/ Ultrasonic NDT Lahivaara et al. [244] Wang et al. [135] Rao et al. [6] Hong et al. [245] Prakash et al. [246] Song et al. [247] Siljama et al. [248] Rudolph et al. [136] Sen et al. [249] Jiang et al. [250] Bouzenad et al. [251] Akcay et al. [252] Luiken et al. [138] Shukla et al. [137] Virupakshappa et al. [253] Kraljevski et al. [254] Wang et al. [255] Rautela et al. [256] Sawant et al. [257]

### v2

@@ Appendix D Open-Source Software and Data Availability
  • Medical Ultrasound Tomography. In Refs. [76, 225], the authors develop a three-dimensional anatomically realistic numerical phantom of the female breast and a high-fidelity three-dimensional imaging model of ring-array USCT breast ultrasound. The goal of the generated numerical breast phantoms (NBPs) and corresponding simulated measurements is to provide a dataset to 1) support the development and assessment (using objective measurement of image quality) of model-based and learned image reconstruction methods and 2) conduct virtual imaging trials to guide the design of new USCT breast imaging systems. Compared to other acoustic NBPs in the literature, the ones in Ref. [76] are particularly relevant to the development of new learning-based USCT image reconstruction method since they 1) comprise realistic structures and acoustic properties; 2) include lesions and/or other pathologies; and 3) are representative of the stochastic variability in breast size, shape, composition, anatomy, and tissue properties observed in a specified cohort of to-be-imaged subjects. Leveraging tools developed and released by the Virtual Imaging Clinical Trials for Regulatory Evaluation (VICTRE) project [226, 227] of the Food and Drug Administration (FDA), the NBPs in Ref. [76] correspond to the four different levels of breast density defined according to the American College of Radiology’s (ACR) Breast Imaging Reporting and Data System (BI-RADS) [228]: A) Breast is almost entirely fat, B) Breast has scattered areas of fibroglandular density, C) Breast is heterogeneously dense, and D) Breast is extremely dense. Each NBP is a 3D voxelized map consisting of ten tissue types: fat, skin, glandular, nipple, ligament (connective tissue), muscle, terminal duct lobular unit, duct, artery, and vein. In the framework of the authors, large ensembles of stochastic NBPs with realistic variability in breast volume, shape, fraction of glandular tissue, ligament orientation, tissue anatomy, can be generated by controlling input parameters and selecting the random seed number. A variety of lesions (e.g., circumscribed or spiculated masses) can also be inserted at physiologically plausible locations. Tissue-specific acoustic properties (density, SOS, and AA) are then stochastically assigned to each tissue structure within physiological ranges and acoustic heterogeneity within fatty and glandular tissues is also modeled as spatially correlated random fields. Realization of these numerical phantoms and corresponding simulated measurements can be downloaded from Refs. [229, 230, 231], and additional NBPs can be generated using the open-source software in Ref. [232]. [233] presents a large-scale, anatomically realistic ultrasound dataset designed to benchmark neural solvers for wave equations. Featuring over 16 million simulations across diverse breast phantoms, it enables evaluation of neural operators in both forward and inverse tasks under clinically relevant conditions. The dataset addresses scalability, realism, and generalization gaps, and is publicly available to support reproducible research in neural wave-based imaging. Last but not least, [234] introduces the first large-scale open dataset specifically designed for limited-view prostate USCT. It provides over 280,000 paired samples of realistic 2D speed-of-sound phantoms and corresponding full-waveform ultrasound data, derived from anatomically accurate 3D prostate models based on clinical MRI/CT and expert annotations. This dataset addresses a critical gap in prostate cancer imaging research, enabling development and benchmarking of neural inversion methods under realistic limited-angle conditions. Baseline experiments show that while deep learning significantly outperforms traditional physics-based methods in both speed and accuracy, there remains room for improvement before achieving clinical-grade resolution.
TABLE IV: Paper Summary
@@
Reference Pure ML Methods Hybrid Methods Supervised Semi-Supervised Self-Supervised Unsupervised/ Unlabeled Simulations/ Measurements-Driven Physics Aware Parameterized Solution Plug & Play Seismic Wu and Lin [51] Araya-Polo et al. [1] Renan et al. [58] Wang et al. [235] Kazei et al. [236] Wu et al. [237] Sun et al. [56] Hu et al. [238] Zhu et al. [92] Saad et al. [62] Wu et al. [239] Fu et al. [110] Vamaraju et al. [240] Rasht-Behesh et al. [84] Izzatullah et al. [111] Zhang and Wang [95] Medical Ultrasound Lozenski et al. [9] Vedula et al. [241] Feigin et al. [242] Zhao et al. [243] Tong et al. [244] Liu et al. [245] Zhang et al. [246] Zhang et al. [247] Dai et al. [248] Dai et al. [249] Fan et al. [250] Acoustic/ Ultrasonic NDT Lahivaara et al. [251] Wang et al. [152] Rao et al. [6] Hong et al. [252] Prakash et al. [253] Song et al. [254] Siljama et al. [255] Rudolph et al. [153] Sen et al. [256] Jiang et al. [257] Bouzenad et al. [258] Akcay et al. [259] Luiken et al. [155] Shukla et al. [154] Virupakshappa et al. [260] Kraljevski et al. [261] Wang et al. [262] Rautela et al. [263] Sawant et al. [264]
Appendix E List of Acronyms
For the convenience of the reader, this appendix provides a list of acronyms used throughout the manuscript. Each term is defined below to ensure clarity and consistency.
ADMM
Alternating Direction Method of Multipliers
AE
Autoencoder
AI
Artificial Intelligence
CNN
Convolutional Neural Network
CT
Computed Tomography
CWI
Computational Wave Imaging
EFWI
Elastic Full Waveform Inversion
FCN
Fully Convolutional Network
FNO
Fourier Neural Operator
FWI
Full Waveform Inversion
FWIGAN
Full Waveform Inversion Generative Adversarial Network
GAN
Generative Adversarial Network
KL
Kullback-Leibler
MAE
Mean Absolute Error
MAP
Maximum A Posteriori
MCMC
Markov Chain Monte Carlo
ML
Machine Learning
ML-NDT
Machine Learning for Nondestructive Testing
MRI
Magnetic Resonance Imaging
MSE
Mean Squared Error
NDT
Nondestructive Testing
OOD
Out of Distribution
PDE
Partial Differential Equation
PINN
Physics-Informed Neural Network
PPP
Plug-and-Play Prior
PPP-CWI
Plug-and-Play Prior for Computational Wave Imaging
SDE
Stochastic Differential Equation
SNR
Signal-to-Noise Ratio
SOS
Speed of Sound
SSIM
Structural Similarity Index
SSL
Self-Supervised Learning
TV
Total Variation
UPFWI
Unsupervised Physics-Informed Full Waveform Inversion
UQ
Uncertainty Quantification
USCT
Ultrasound Computed Tomography
VAE
Variational Autoencoder